. Determine the y-intercept and the vertex of the of the parabola y = 2x^2+8x+12. Are there any x-intercepts.
Answer by stanbon(75887)

Determine the y-intercept and the vertex of the of the parabola
y = 2x^2+8x+12. Are there any x-intercepts.
----------------
y-int:: f(0) = 2*0+8*0+12 = 12
-------------------------------------
Vertex::
The vertex occurs when x = -b/(2a) = -8/(2*2) = -2
f(-2) = 2(-2)^2+8(-2)+12 = 8 - 16 + 12 = 4
Vertex:: (-2,4)
--------------------
Cheers,
Stan H.
-----------------